BigPeen:

Five added to twice Erik’s age is the same as 3 times his age minus 2. How old is Erik?

xxDeppressionxx:

Welcome to Qc!, Do you happen to have any answer choices?

jhonyy9:

Erik's age let be x 5 +2x = 3x -2 solve this for x

xxDeppressionxx:

ok then so the first thing that you have to do is multiply 7 and 2 then you do that minus 3 which eaquals 7

jhonyy9:

@Laylalyssa any idea here pls ? ty

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxDeppressionxx ok then so the first thing that you have to do is multiply 7 and 2 then you do that minus 3 which eaquals 7 \(\color{#0cbb34}{\text{End of Quote}}\) but u didnt do subtraction right lol 7 x 2= 14 14-3= 11 how did u get 7 bro?

xxDeppressionxx:

i meant 11

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 Erik's age let be x 5 +2x = 3x -2 solve this for x \(\color{#0cbb34}{\text{End of Quote}}\) just solve this equation for x

xxDeppressionxx:

7 • 2 = 14 - 3 = 11

BigPeen:

alright ;)

jhonyy9:

Five added to twice Erik’s age is the same as 3 times his age minus 2. how you get this 7 ? from where ?

jhonyy9:

Erik's age note an x twice Erik's age is 2x five added to twice Erik's age is 5 +2x 3 times his age mean 3x minus 2 is 3x-2 so we have 5 +2x = 3x -2 solve this for x

jhonyy9:

questions ? this is so easy !!! just you need solve this equation for x ...